# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 2.3. Filtered symplectic homology
In fact, this is also true for the symplectic homology over any ring.
@@ 2.4.1. Basic definitions and facts
Let $x$ be an isolated closed Reeb orbit, not necessarily prime. Fix an isolating neighborhood $U$ of $x({\mathbb{R}})$ in $M$. (By definition, $U$ is a neighborhood of $x$ containing no other closed orbit of period close to ${\mathcal{A}}(x)$.) Consider the product $\widehat{U}=U\times(1-\epsilon,1+\epsilon)\subset\widehat{W}$ and let $H=h(r)$ be a Hamiltonian such that $h^{\prime}(1)={\mathcal{A}}(x)$ and $h^{\prime\prime}(1)>0$. Then $x$ gives rise to an isolated set $S$ of 1-periodic orbits of $H$ with initial conditions on $x({\mathbb{R}})$. By definition, the symplectic homology $\operatorname{SH}(x;{\mathbb{F}})$ is the local Floer homology HF($\operatorname{HF}(H;{\mathbb{F}})$ of $H$ at $S$, and $\operatorname{CH}(x;{\mathbb{F}})$ is the $S^{1}$-equivariant local Floer homology.
Specifically, let $F$ be a non-degenerate 1-periodic in time perturbation of $H$ and $J$ a 1-periodic in time almost complex structure. When $F$ is sufficiently $C^{2}$-close to $H$, all 1-periodic orbits of the flow of $F$ that $x$ splits into and Floer cylinders between them are contained in $\widehat{U}$. (This a variant of Gromov’s compactness, which follows, for instance, from the estimates in [Sa99, Sec. 1.5].) For a generic choice of $J$ the regularity conditions are satisfied. Hence we obtain a complex and $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of this complex. The equivariant version is defined in a similar fashion with appropriate modifications; see [GG20, Sec. 2.3].
@@
Similarly to the filtered homology, $\operatorname{SH}(x;{\mathbb{F}})$ and $\operatorname{CH}(x;{\mathbb{Q}})$ can be more explicitly described as follows. Consider, a small non-degenerate perturbation of $\alpha$. Then $x$ splits into a finite collection of non-degenerate orbits $x_{i}$. Essentially by definition, $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of a certain complex with generators $\check{x}_{i}$ and $\hat{x}_{i}$ of degrees $\mu(x_{i})$ and, respectively, $\mu(x_{i})+1$ and the differential coming from the Floer–Morse–Bott construction as in [Bo, BO09a] and also, e.g., [ÇG2M23]. Likewise, $\operatorname{CH}(x;{\mathbb{Q}})$ is the homology of a complex over ${\mathbb{Q}}$ generated by the orbits $x_{i}$ and graded by the Conley–Zehnder index; see [GG20, Sec. 3]. This is no longer literally true when ${\mathbb{Q}}$ is replaced by a field of positive characteristic; see Remark 2.6.
@@
Moreover, by [GG20, Prop 2.21], the shift operator ${\mathcal{D}}$ vanishes in local symplectic homology when ${\mathbb{F}}={\mathbb{Q}}$, and the Gysin sequence splits into short exact sequences. (This is no longer true when ${\mathit{c}har}\,{\mathbb{F}}>0$ unless ${\mathit{c}har}\,{\mathbb{F}}$ and the order of iteration of $x$ are relatively prime.) As a consequence,
@@
As was mentioned above, a closed orbit $x$ is said to be ${\mathbb{F}}$-visible if $\operatorname{SH}(x;{\mathbb{F}})\neq 0$. Likewise, $x$ is equivariantly ${\mathbb{F}}$-visible if $\operatorname{CH}(x;{\mathbb{F}})\neq 0$. An ${\mathbb{F}}$-visible closed orbit is automatically equivariantly ${\mathbb{F}}$-visible by the Gysin sequence. The converse is also true when ${\mathbb{F}}={\mathbb{Q}}$ due to (2.10).
@@
$\chi(x;{\mathbb{F}}):=\sum_{m}(-1)^{m}\dim\operatorname{SH}_{m}(x;{\mathbb{F}})=0.$
Therefore,
$\dim\operatorname{SH}(x;{\mathbb{F}})\textrm{ is always even,}$,
@@
[…]
@@
For instance, when $x$ is non-degenerate, $\chi^{\scriptscriptstyle{eq}}(x)$ is either $(-1)^{\mu(x)}$ or 00 depending on whether $x$ is good or bad. We note that here we restrict our background field ${\mathbb{F}}$ to ${\mathbb{Q}}$ because otherwise $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ can be infinite and then the right-hand side of (2.12) is not defined.
It readily follows from the description of $\operatorname{CH}(x;{\mathbb{Q}})$ in Section 2.4.1 that, whenever $x$ is prime, $\chi^{\scriptscriptstyle{eq}}(x)$ is equal to the Hopf index of the Poincaré return map of $x$ up to the sign $(-1)^{n}$. This is no longer true when $x$ is iterated. For instance, $\chi^{\scriptscriptstyle{eq}}(x)=0$ when $x$ is non-degenerate and bad, while the Hopf index is ±$\pm 1$. Moreover, let $x=z^{k}$ where $z$ is prime and, for the sake of simplicity, $k$ is odd. Then $\chi^{\scriptscriptstyle{eq}}(x)$ counts the number (with signs) of $k$-periodic orbits of a small non-degenerate perturbation of the return map $\varphi$ of $z$, while the Hopf index of $\varphi^{k}$ counts (with the same signs) the number of $k$-periodic points.
@@ 2.4.2. Behavior under iterations
Recall that $k\in{\mathbb{N}}$ or an iterate $x^{k}$ is said to be admissible for $x$ if $k$ is not divisible by the degree of any root of unity among the eigenvalues of the linearized Poincaré return map of $x$. For instance, all $k\in{\mathbb{N}}$ are admissible when $x$ is totally degenerate, i.e., all eigenvalues are equal to 1.
@@
Let $x$ be an isolated (not necessarily prime) closed Reeb orbit. Assume that $k$ is admissible for $x$ and odd if the number of eigenvalues of the return map of $x$ in $(-1,0)$ is odd. Then
@@
We claim that $x_{i}^{k}$ is good (bad) if and only if so is $x_{i}$, and that $\mu(x_{i})$ and $\mu\big{(}x^{k}_{i}\big{)}$ have the same parity. When $k$ is odd this is obvious. Indeed, the eigenvalues in $(-1,0)$ remain in that range for all iterates and additional eigenvalues come in pairs, which implies the parity statement. Likewise, an odd iterate of a good (bad) closed orbit is automatically good (bad). If $-1$ is an eigenvalue of the return map of $x$, the iteration $k$ is necessarily odd, since $k$ is admissible, and the claim follows. Assume that $-1$ is not an eigenvalue and $k$ is even. Then for each $x_{i}$ the parity of the number of real eigenvalues in $(-1,0)$ is the same as for $x$ and, by the conditions of the lemma, even. Hence, in this case, $\mu(x_{i})$ and $\mu\big{(}x^{k}_{i}\big{)}$ also have the same parity. Furthermore, if $x_{i}=z^{r}$ where $z$ is prime, $x_{i}$ is good (bad) if and only $\mu(x_{i})$ and $\mu(z)$ have the same (opposite) parity. Ditto, for $x_{i}^{k}=z^{kr}$. As a consequence $x_{i}$ and $x_{i}^{k}$ are good (bad) simultaneously.
@@
Note that here the orbit $x$ need not be prime and $k$ is not required to be admissible. However, $x^{k}$ must be isolated. Furthermore, the condition that ${\mathbb{F}}={\mathbb{Q}}$ in (2.14) can be relaxed. It would be sufficient to assume that ${\mathit{c}har}\,{\mathbb{F}}$ and $k$ are relatively prime.
@@
In the notation from Section 2.4.1, let $H=h(r)\colon\widehat{U}\to{\mathbb{R}}$ be a Hamiltonian from the construction of $\operatorname{SH}\big{(}x^{k};{\mathbb{F}}\big{)}:=\operatorname{HF}(H;{\mathbb{F}})$. Fix a non-degenerate 1-periodic in time perturbation $F$ of $H$ and a regular, also 1-periodic in time, almost complex structure $J$ on $\widehat{U}$. Thus, in particular, $\operatorname{HF}(H;{\mathbb{F}})=\operatorname{HF}(F;{\mathbb{F}})$. Denote by $H_{k}$, $F_{k}$ and $J_{k}$ the pull-backs of the Hamiltonians and $J$ to the $k$-fold covering $\pi\colon\widehat{U}_{k}\to\widehat{U}$. Then $\operatorname{SH}(kx;{\mathbb{F}}):=\operatorname{HF}(H_{k};{\mathbb{F}})=\operatorname{HF}(F_{k};{\mathbb{F}})$. These pull-backs are invariant under the ${\mathbb{Z}}_{k}$-action on $\widehat{U}_{k}$ by deck transformations. The 1-periodic orbits of $F_{k}$ project to the periodic orbits of $F$, and so do Floer cylinders. Moreover, on the sets of 1-periodic orbits and Floer cylinders these projections are onto and $k$-to-1. To be more precise, since all 1-periodic orbits of $F$ are in the free homotopy class $k[S^{1}]$, the Hamiltonian $F_{k}$ has exactly $k$ distinct 1-periodic orbits lying above every 1-periodic orbit of $F$, and deck transformations act freely on these orbits. It follows that the same is true for Floer cylinders; cf. [Sa99, Prop. 1.21].
We claim that the pair $(F_{k},J_{k})$ is regular if and only if the pair $(F,J)$ is regular; cf. [KS, Prop. 5.13]. Indeed, let $u$ be a Floer cylinder in $\widehat{U}_{k}$. Then the pull-backs $u^{*}T\widehat{U}_{k}$ and $(\pi u)^{*}T\widehat{U}$ are naturally isomorphic. (This is true whenever $\pi$ is a local diffeomorphism.) Hence, we obtain an isomorphism between the Sobolev spaces of sections of these bundles; see [Sa99, Sec. 2] for the definitions. By construction, this isomorphism intertwines the linearized perturbed Cauchy–Riemann operators $D_{u}$ along $u$ and $D_{\pi u}$ along $\pi u$. Therefore, $D_{u}$ is onto if and only if $D_{\pi u}$ is onto. Finally, since every Floer cylinder in $\widehat{U}$ has a lift to $\widehat{U}_{k}$, we conclude that the pairs $(F,J)$ and $(F_{k},J_{k})$ are regular simultaneously.
@@
When ${\mathbb{F}}={\mathbb{Q}}$, the Floer complex $\operatorname{CF}(F,J)$ is naturally isomorphic to the invariant part $\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}$ of $\operatorname{CF}(H_{k},J_{k})$. Indeed, in one direction we have (for any ${\mathbb{F}})$ the homomorphism of complexes $\operatorname{CF}(H_{k},J_{k})\to\operatorname{CF}(F,J)$ induced by the projection $\pi\colon\widehat{U}_{k}\to\widehat{U}$. In the opposite direction, we have the averaging map. Namely, for $z\in\operatorname{CF}(F,J)$, pick an arbitrary lift $z^{\prime}\in\operatorname{CF}(H_{k},J_{k})$ and set, in the obvious notation,
@@
This map gives rise to an isomorphism $\operatorname{CF}(H,J)\cong\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}$, which concludes the proof of (2.14). ∎
@@
The behavior under iterations of local symplectic or Floer homology (equivariant or not) is complicated and overall poorly understood. For instance, it is not known if a closed orbit can go from invisible to visible under iterations, i.e., if a closed orbit $x^{k}$ can be ${\mathbb{F}}$-visible when $x$ is prime and ${\mathbb{F}}$-invisible, without additional assumptions on $x$, the ground field ${\mathbb{F}}$ or $k$. At the same time, $x^{k}$ can in general become ${\mathbb{F}}$-invisible when $x$ is prime and ${\mathbb{F}}$-visible. A simple situation when this happens is when $x$ is non-degenerate and alternating, $k$ is even and ${\mathbb{F}}={\mathbb{Q}}$. In Examples 2.12 and 2.13 we describe more interesting cases of this phenomenon. Lemma 2.8 provides a simple non-vanishing criterion, which is sufficient for our purposes. (The question for a variant of local symplectic homology and admissible iterations has been investigated in [HHM]. The homology there is defined via generating functions and is hypothetically isomorphic to the local homology used here; cf. Remark 4.3.)
@@
Let $\psi\colon{\mathbb{R}}^{2}={\mathbb{C}}\to{\mathbb{C}}={\mathbb{R}}^{2}$ be the time-one flow of the degree-three monkey saddle Hamiltonian $H=\Re(z^{3})$ and let $R_{2\pi/3}$ be the rotation of ${\mathbb{R}}^{2}$ counterclockwise by $32\pi/$ 3. (We view $R_{2\pi/3}$ as an element of the universal covering $\operatorname{\widetilde{Sp}}(2)$.) Note that $R_{2\pi/3}$ and the flow of $H$ commute. Set $\varphi=R_{2\pi/3}\psi$ and let $x$ be a closed Reeb orbit with the return map $\varphi$. The orbit $x$ is non-degenerate, $\operatorname{CH}(x;{\mathbb{Q}})={\mathbb{Q}}$ and for a suitable trivialization $\operatorname{CH}(x;{\mathbb{Q}})$ is supported in degree one. The return map of $x^{3}$ is $\varphi^{3}$, which is the time-three flow of $H$. One can show that $\operatorname{CH}(x^{3};{\mathbb{Q}})=0$. (This is not obvious.) However, $\operatorname{HF}(\varphi^{3};{\mathbb{Q}})$ is ${\mathbb{Q}}^{2}$ and supported in degree zero. This gives another example where the inequality in (2.16) is strict; see also Example 2.13 below.
@@
The notion of an (non-)alternating closed orbit is not particularly useful in the setting where $x$ is not strongly non-degenerate and the return map $\varphi$ has roots of unity of even degree among its eigenvalues. The reason is that in this case the behavior of the local symplectic homology is not determined by the eigenvalues and depends on higher order terms. For instance, in the spirit of Example 2.12, let $R_{\pi}\colon{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be the counterclockwise rotation by $\pi$, i.e., central symmetry, viewed as an element of $\operatorname{\widetilde{Sp}}(2)$. Consider the Hamiltonians $H_{0}=p^{4}-q^{4}$ and $H_{1}=(p^{2}+q^{2})^{2}$ on ${\mathbb{R}}^{2}$, where $(p,q)$ are the Darboux coordinates. The flows of $H_{0}$ and $H_{1}$ commute with $R_{\pi}$. Let $\varphi_{i}$, for $i=0,1$, be obtained by composing $R_{\pi}$ with the time-one flow of $H_{i}$ and $x_{i}$ be a closed Reeb orbit with the return map $\varphi_{i}$. Clearly, $\operatorname{CH}(x_{i};{\mathbb{Q}})=\operatorname{HF}(\varphi_{i})={\mathbb{Q}}$ is supported in degree one. However, one can show that $\operatorname{CH}\big{(}x_{0}^{2};{\mathbb{Q}}\big{)}=0$ and $\operatorname{HF}\big{(}\varphi_{0}^{2};{\mathbb{Q}}\big{)}={\mathbb{Q}}$ (supported in degree two). At the same time, $\operatorname{CH}\big{(}x_{1}^{2};{\mathbb{Q}}\big{)}=\operatorname{HF}\big{(}\varphi_{1}^{2};{\mathbb{Q}}\big{)}={\mathbb{Q}}$ (supported in degree three).

### v3

@@ 2.3. Filtered symplectic homology
While the construction of symplectic homology in [BO13] uses $\mathbb{Q}$} as the coefficient field and a semi-infinite interval as $I$, it carries over word-for-word to $\mathbb{Z}$}, and hence any ring, and an arbitrary interval $I$; see, e.g., [GG20, Sect. 2] for a detailed discussion of $\mathcal{D}$} and [Abu] for a non-equivariant construction over $\mathbb{Z}$}.
@@ 2.4.1. Basic definitions and facts
Let $x$ be an isolated closed Reeb orbit, not necessarily prime. Fix an isolating neighborhood $U$ of $x({\mathbb{R}})$ in $M$. (By definition, $U$ is a neighborhood of $x$ containing no other closed orbit of period close to ${\mathcal{A}}(x)$.) Consider the product $\widehat{U}=U\times(1-\epsilon,1+\epsilon)\subset\widehat{W}$ and let $H=h(r)$ be a Hamiltonian such that $h^{\prime}(1)={\mathcal{A}}(x)$ and $h^{\prime\prime}(1)>0$. Then $x$ gives rise to an isolated set $S$ of 1-periodic orbits of $H$ with initial conditions on $x({\mathbb{R}})$. By definition, the symplectic homology $\operatorname{SH}(x;{\mathbb{F}})$ is the local Floer homology $\operatorname{HF}(H;{\mathbb{F}})$ of $H$ at $S$, and $\operatorname{CH}(x;{\mathbb{F}})$ is the $S^{1}$-equivariant local Floer homology.
Specifically, let $F$ be a non-degenerate 1-periodic in time perturbation of $H$ and $J$ a 1-periodic in time almost complex structure. When $F$ is sufficiently $C^{2}$-close to $H$, all 1-periodic orbits of the flow of $F$ that $x$ splits into and Floer cylinders between them are contained in $\widehat{U}$. (This is a variant of Gromov’s compactness, which follows, for instance, from the estimates in [Sa, Sec. 1.5].) For a generic choice of $J$ the regularity conditions are satisfied. Hence we obtain a complex and $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of this complex. The equivariant version is defined in a similar fashion with appropriate modifications; see [GG20, Sec. 2.3].
@@
Similarly to the filtered homology, $\operatorname{SH}(x;{\mathbb{F}})$ and $\operatorname{CH}(x;{\mathbb{Q}})$ can be more explicitly described as follows. Consider a small non-degenerate perturbation of $\alpha$. Then $x$ splits into a finite collection of non-degenerate orbits $x_{i}$. Essentially by definition, $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of a certain complex with generators $\check{x}_{i}$ and $\hat{x}_{i}$ of degrees $\mu(x_{i})$ and, respectively, $\mu(x_{i})+1$ and the differential coming from the Floer–Morse–Bott construction as in [Bo, BO09a] and also, e.g., [ÇG2M23]. Likewise, $\operatorname{CH}(x;{\mathbb{Q}})$ is the homology of a complex over $\mathbb{Q}$} generated by the orbits $x_{i}$ and graded by the Conley–Zehnder index; see [GG20, Sec. 3]. This is no longer literally true when $\mathbb{Q}$} is replaced by a field of positive characteristic; see Remark 2.6.
@@
This sequence is just a local variant of the Gysin sequence from [BO13]. Moreover, by [GG20, Cor. 3.7], the shift operator $\mathcal{D}$} vanishes in local symplectic homology when $\mathbb{F}$$\mathbb{Q}$}, and the Gysin sequence splits into short exact sequences. (This is no longer true when ${\mathit{c}har}\,{\mathbb{F}}>0$ unless ${\mathit{c}har}\,{\mathbb{F}}$ and the order of iteration of $x$ are relatively prime.) As a consequence,
@@
As was mentioned above, a closed orbit $x$ is said to be $\mathbb{F}$}-visible if $\operatorname{SH}(x;{\mathbb{F}})\neq 0$. Likewise, $x$ is equivariantly $\mathbb{F}$}-visible if $\operatorname{CH}(x;{\mathbb{F}})\neq 0$. An $\mathbb{F}$}-visible closed orbit is automatically equivariantly $\mathbb{F}$}-visible by the Gysin sequence. The converse is also true when $\mathbb{F}$$\mathbb{Q}$} due to (2.10).
@@
$\chi(x;{\mathbb{F}}):=\sum_{m}(-1)^{m}\dim\operatorname{SH}_{m}(x;{\mathbb{F}})=0.$ (2.11)
Indeed, as stated in the beginning of this section $\operatorname{SH}(x;{\mathbb{F}})$ is the homology of a complex with generators coming in pairs $\check{x}_{i}$ and $\hat{x}_{i}$ where the degree of $\check{x}_{i}$ is $\mu(x_{i})$ and the degree of $\hat{x}_{i}$ is $\mu(x_{i})+1$. Therefore,
$\dim\operatorname{SH}(x;{\mathbb{F}})\textrm{ is always even,}$
@@
$\dim\operatorname{SH}(x;{\mathbb{F}})\geq 2\textrm{ when $x$ is ${\mathbb{F}}$-visible.}$ (2.12)
@@
For instance, when $x$ is non-degenerate, $\chi^{\scriptscriptstyle{eq}}(x)$ is either $(-1)^{\mu(x)}$ or 0 depending on whether $x$ is good or bad. We note that here we restrict our background field $\mathbb{F}$} to $\mathbb{Q}$} because otherwise $\operatorname{supp}\operatorname{CH}(x;{\mathbb{F}})$ can be infinite and then the right-hand side of (2.13) is not defined.
It readily follows from the description of $\operatorname{CH}(x;{\mathbb{Q}})$ in Section 2.4.1 that, whenever $x$ is prime, $\chi^{\scriptscriptstyle{eq}}(x)$ is equal to the Hopf index of the Poincaré return map of $x$ up to the sign $(-1)^{n}$. This is no longer true when $x$ is iterated. For instance, $\chi^{\scriptscriptstyle{eq}}(x)=0$ when $x$ is non-degenerate and bad, while the Hopf index is $\pm 1$. Moreover, let $x=z^{k}$ where $z$ is prime and, for the sake of simplicity, $k$ is odd. Then $\chi^{\scriptscriptstyle{eq}}(x)$ counts the number (with signs) of $k$-periodic orbits of a small non-degenerate perturbation of the return map $\varphi$ of $z$, while the Hopf index of $\varphi^{k}$ counts (with the same signs) the number of $k$-periodic points.
@@ 2.4.2. Behavior under iterations
Recall that a positive integer $k$ and the iterate $x^{k}$ are said to be admissible for $x$ if $k$ is not divisible by the degree of any root of unity other than $1$ among the eigenvalues of the linearized Poincaré return map of $x$. For instance, all $k\in{\mathbb{N}}$ are admissible when $x$ is totally degenerate, i.e., all eigenvalues are equal to 1.
@@
Let $x$ be an isolated (not necessarily prime) closed Reeb orbit. Assume that $k$ is admissible for $x$ and odd if the number of eigenvalues of the return map of $x$ in (1,0)(-1,0) is odd. Then
@@
We claim that $x_{i}^{k}$ is good (bad) if and only if so is $x_{i}$, and that $\mu(x_{i})$ and $\mu\big(x^{k}_{i}\big)$ have the same parity. When $k$ is odd this is obvious. Indeed, the eigenvalues in (1,0)(-1,0) remain in that range for all iterates and additional eigenvalues come in pairs, which implies the parity statement. Likewise, an odd iterate of a good (bad) closed orbit is automatically good (bad). If 1-1 is an eigenvalue of the return map of $x$, the iteration $k$ is necessarily odd, since $k$ is admissible, and the claim follows. Assume that 1-1 is not an eigenvalue and $k$ is even. Then for each $x_{i}$ the parity of the number of real eigenvalues in (1,0)(-1,0) is the same as for $x$ and, by the conditions of the lemma, even. Hence, in this case, $\mu(x_{i})$ and $\mu\big(x^{k}_{i}\big)$ also have the same parity. Furthermore, if $x_{i}=z^{r}$ where $z$ is prime, $x_{i}$ is good (bad) if and only $\mu(x_{i})$ and $\mu(z)$ have the same (opposite) parity. This is also true for $x_{i}^{k}=z^{kr}$. As a consequence $x_{i}$ and $x_{i}^{k}$ are good (bad) simultaneously.
@@
Note that here the orbit $x$ need not be prime and $k$ is not required to be admissible. However, $x^{k}$ must be isolated. Furthermore, the condition that $\mathbb{F}$$\mathbb{Q}$} in (2.15) can be relaxed. It would be sufficient to assume that ${\mathit{c}har}\,{\mathbb{F}}$ and $k$ are relatively prime.
@@
In the notation from Section 2.4.1, let $H=h(r)\colon\widehat{U}\to{\mathbb{R}}$ be a Hamiltonian from the construction of $\operatorname{SH}\big(x^{k};{\mathbb{F}}\big):=\operatorname{HF}(H;{\mathbb{F}})$. Fix a non-degenerate 1-periodic in time perturbation $F$ of $H$ and a regular 1-periodic almost complex structure $J$ on $\widehat{U}$. Thus, in particular, $\operatorname{HF}(H;{\mathbb{F}})=\operatorname{HF}(F;{\mathbb{F}})$. Denote by $H_{k}$, $F_{k}$ and $J_{k}$ the pull-backs of the Hamiltonians and $J$ to the $k$-fold covering $\pi\colon\widehat{U}_{k}\to\widehat{U}$. Then $\operatorname{SH}(kx;{\mathbb{F}}):=\operatorname{HF}(H_{k};{\mathbb{F}})=\operatorname{HF}(F_{k};{\mathbb{F}})$. These pull-backs are invariant under the ${\mathbb{Z}}_{k}$-action on $\widehat{U}_{k}$ by deck transformations. The 1-periodic orbits of $F_{k}$ project to the periodic orbits of $F$, and so do Floer cylinders. Moreover, on the sets of 1-periodic orbits and Floer cylinders these projections are onto and $k$-to-1. To be more precise, since all 1-periodic orbits of $F$ are in the free homotopy class $k[S^{1}]$, the Hamiltonian $F_{k}$ has exactly $k$ distinct 1-periodic orbits lying above every 1-periodic orbit of $F$, and deck transformations act freely on these orbits. It follows that the same is true for Floer cylinders; cf. [Sa, Prop. 1.21].
We claim that the pair $(F_{k},J_{k})$ is regular if and only if the pair $(F,J)$ is regular; cf. [KS, Prop. 5.13]. Indeed, let $u$ be a Floer cylinder in $\widehat{U}_{k}$. Then the pull-backs $u^{*}T\widehat{U}_{k}$ and $(\pi u)^{*}T\widehat{U}$ are naturally isomorphic. (This is true whenever $\pi$ is a local diffeomorphism.) Hence, we obtain an isomorphism between the Sobolev spaces of sections of these bundles; see [Sa, Sec. 2] for the definitions. By construction, this isomorphism intertwines the linearized perturbed Cauchy–Riemann operators $D_{u}$ along $u$ and $D_{\pi u}$ along $\pi u$. Therefore, $D_{u}$ is onto if and only if $D_{\pi u}$ is onto. Finally, since every Floer cylinder in $\widehat{U}$ has a lift to $\widehat{U}_{k}$, we conclude that the pairs $(F,J)$ and $(F_{k},J_{k})$ are regular simultaneously.
@@
When $\mathbb{F}$$\mathbb{Q}$}, the Floer complex $\operatorname{CF}(F,J)$ is naturally isomorphic to the invariant part $\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}$ of $\operatorname{CF}(H_{k},J_{k})$. Indeed, in one direction we have (for any ${\mathbb{F}})$ the homomorphism of complexes $\operatorname{CF}(H_{k},J_{k})\to\operatorname{CF}(F,J)$ induced by the projection $\pi\colon\widehat{U}_{k}\to\widehat{U}$. In the opposite direction, we have the averaging map. Namely, for $z\in\operatorname{CF}(F,J)$, pick an arbitrary lift $z^{\prime}\in\operatorname{CF}(H_{k},J_{k})$ and set, in the obvious notation,
@@
This map gives rise to an isomorphism $\operatorname{CF}(F,J)\cong\operatorname{CF}(H_{k},J_{k})^{{\mathbb{Z}}_{k}}$, which concludes the proof of (2.15). ∎
@@
The behavior under iterations of local symplectic or Floer homology (equivariant or not) is complicated and overall poorly understood. For instance, it is not known if a closed orbit can go from invisible to visible under iterations, i.e., if a closed orbit $x^{k}$ can be $\mathbb{F}$}-visible when $x$ is prime and $\mathbb{F}$}-invisible, without additional assumptions on $x$, the ground field $\mathbb{F}$} or $k$. At the same time, $x^{k}$ can in general become $\mathbb{F}$}-invisible when $x$ is prime and $\mathbb{F}$}-visible. A simple situation when this happens is when $x$ is non-degenerate and alternating, $k$ is even and $\mathbb{F}$$\mathbb{Q}$}. In Examples 2.12 and 2.13 we describe more interesting cases of this phenomenon. Lemma 2.8 provides a simple non-vanishing criterion, which is sufficient for our purposes. (The question for a variant of local symplectic homology and admissible iterations has been investigated in [HHM]. The homology there is defined via generating functions and is hypothetically isomorphic to the local homology used here; cf. Remark 4.5.)
@@
Let $\psi\colon{\mathbb{R}}^{2}={\mathbb{C}}\to{\mathbb{C}}={\mathbb{R}}^{2}$ be the time-one flow of the degree-three monkey saddle Hamiltonian $H=\Re(z^{3})$ and let $R_{2\pi/3}$ be the rotation of ${\mathbb{R}}^{2}$ counterclockwise by $2\pi/3$. (We view $R_{2\pi/3}$ as an element of the universal covering $\operatorname{\widetilde{Sp}}(2)$.) Note that $R_{2\pi/3}$ and the flow of $H$ commute. Set $\varphi=R_{2\pi/3}\psi$ and let $x$ be a closed Reeb orbit with the return map $\varphi$. The orbit $x$ is non-degenerate, $\operatorname{CH}(x;{\mathbb{Q}})={\mathbb{Q}}$ and for a suitable trivialization $\operatorname{CH}(x;{\mathbb{Q}})$ is supported in degree one. The return map of $x^{3}$ is $\varphi^{3}$, which is the time-three flow of $H$. One can show that $\operatorname{CH}(x^{3};{\mathbb{Q}})=0$. (This is not obvious.) However, $\operatorname{HF}(\varphi^{3};{\mathbb{Q}})$ is ${\mathbb{Q}}^{2}$ and supported in degree zero. This gives another example where the inequality in (2.17) is strict; see also Example 2.13 below.
@@
The notion of a (non-)alternating closed orbit is not particularly useful in the setting where $x$ is not strongly non-degenerate and the return map $\varphi$ has roots of unity of even degree among its eigenvalues. The reason is that in this case the behavior of the local symplectic homology is not determined by the eigenvalues and depends on higher order terms. For instance, in the spirit of Example 2.12, let $R_{\pi}\colon{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be the counterclockwise rotation by $\pi$, i.e., central symmetry, viewed as an element of $\operatorname{\widetilde{Sp}}(2)$. Consider the Hamiltonians $H_{0}=p^{4}-q^{4}$ and $H_{1}=(p^{2}+q^{2})^{2}$ on ${\mathbb{R}}^{2}$, where $(p,q)$ are the Darboux coordinates. The flows of $H_{0}$ and $H_{1}$ commute with $R_{\pi}$. Let $\varphi_{i}$, for $i=0,1$, be obtained by composing $R_{\pi}$ with the time-one flow of $H_{i}$ and $x_{i}$ be a closed Reeb orbit with the return map $\varphi_{i}$. Clearly, $\operatorname{CH}(x_{i};{\mathbb{Q}})=\operatorname{HF}(\varphi_{i})={\mathbb{Q}}$ is supported in degree one. However, one can show that $\operatorname{CH}\big(x_{0}^{2};{\mathbb{Q}}\big)=0$ and $\operatorname{HF}\big(\varphi_{0}^{2};{\mathbb{Q}}\big)={\mathbb{Q}}$ (supported in degree two). At the same time, $\operatorname{CH}\big(x_{1}^{2};{\mathbb{Q}}\big)=\operatorname{HF}\big(\varphi_{1}^{2};{\mathbb{Q}}\big)={\mathbb{Q}}$ (supported in degree three).